OpenStudy (anonymous):

(Will give medal) How can it be that 1^1 = 1 and 1^0 = 1, but 0^1 = 0? Can you explain it?

OpenStudy (ivettef365):

1^1 = 1 x 1 = 1 a^0 = 1 any number raised to 0 = 1 0^1 = 0 x 0 = 0

OpenStudy (cggurumanjunath):